Two Investments totally $49500. They yield an annual income of $4080 a year. One investment yilds 6% a year and the other yields 9% each year. How much is invested at each rate?

Answer by lwsshak3(11628)

Two Investments totally $49500. They yield an annual income of $4080 a year. One investment yilds 6% a year and the other yields 9% each year. How much is invested at each rate?
***
let x= amount invested at 6% rate
49500-x= amount invested at 9% rate
..
.06x+.09(49500-x)=4080
.06x+4455-.09x=4080
.03x=375
x=12,500
49500-x=37,000
How much is invested at each rate?
$12,500 at 6% rate
$37,000 at 9% rate